OpenStudy (anonymous):

Can someone help me solve this radical equation. √(2x-9)+√(2x+6)=3

OpenStudy (luigi0210):

Welcome to Openstudy :)

OpenStudy (anonymous):

\[\begin{align*}\sqrt{2x-9}+\sqrt{2x+6}&=3\\ \sqrt{2x-9}&=3-\sqrt{2x+6}\\ 2x-9&=\left(3-\sqrt{2x+6}\right)^2\\ 2x-9&=9-6\sqrt{2x+6}+(2x+6)\\ -24&=-6\sqrt{2x+6}\\ &\vdots \end{align*}\] Does that help?

OpenStudy (anonymous):

Thank you this really helped.